OpenStudy (anonymous):

How do I get 3x + 5y <= 340 to 10x + 15y?

OpenStudy (anonymous):

i think you are not getting an answer because the question is not clear

OpenStudy (anonymous):

If @satellite73 couldnt answer it on the spot then u know something is wrong with ur question XD

OpenStudy (anonymous):

I'll type the whole question out... but it's long

OpenStudy (anonymous):

The name of your company The type of business (i.e., clothing, electronics, furniture, etc.) Two products your company will make Assign variables to the two products; please use x and y A system of inequalities based on the following information Produce at least 30 of product #1. Produce at least 20 of product #2. Product #1 costs $3 per unit to make. Product #2 $5 per unit to make. The total production cost cannot exceed $340. Graph the system of inequalities This can be done by hand or using a graphing software like GeoGebra. Label the lines, points of intersection, and axis. Show the shaded solution of the three inequalities and explain what it represents. If GeoGebra is used, follow Directions for GeoGebra. If Product #1 earns a profit of $10 per unit and Product #2 earns a profit of $15 per unit, find the combination of Product #1 and Product #2 that will maximize profit. A commercial that could be used to advertise your two products to the general public. You may include a slogan or jingle.

OpenStudy (anonymous):

I have most of it done..

OpenStudy (anonymous):

ok i see what the question is

OpenStudy (anonymous):

you have graphed \[3x + 5y\leq 340\] and \(x>30,y>20\)

OpenStudy (anonymous):

and you want to maximize \(10x+5y\)

OpenStudy (anonymous):

so what you want to do, if i remember correctly, is look at the corners of the graph where the lines intersect then check which one makes \(10x+5y\) the largest

OpenStudy (anonymous):

Did 10x+5y come from 3x+5y? and I must of graphed it wrong?

OpenStudy (anonymous):

no no that was a typo on my part, sorry

OpenStudy (anonymous):

you want to maximize \(10x+15y\) is what i meant to write

OpenStudy (anonymous):

i wrote it twice too, didn't mean to confuse you, sorry about that

OpenStudy (anonymous):

so can you see the corners of your graph? where the lines intersect?

OpenStudy (anonymous):

You're fine. but I still don't understand where 10x+15y came from. The VERY first thing I posted you can ignore!

OpenStudy (anonymous):

it came from this like here in the question "If Product #1 earns a profit of $10 per unit and Product #2 earns a profit of $15 per unit, find the combination of Product #1 and Product #2 that will maximize profit."

OpenStudy (anonymous):

the profit is \(10x+15y\)

OpenStudy (anonymous):

This is my graph.

OpenStudy (anonymous):

OMG. I AM SO SORRY! I am stupid.. I see now..

OpenStudy (anonymous):

i can't open it, but it should look something like this http://www.wolframalpha.com/input/?i=3x%2B5y%3C340%2C+x%3E+30%2C+y%3E20

OpenStudy (anonymous):

you have to check the corners of the graph to see which numbers make \(10x+15y\) the largest

OpenStudy (anonymous):

It look a little like that! But more of just a line.. Not a triangle.

OpenStudy (anonymous):

you have to graph 3 lines the vertical line \(x=30\) and shade to the right also the horizontal line \(y=20\) and shade above it then \(5x+3y=340\) and shade below it

OpenStudy (anonymous):

that is why you get a triangle

OpenStudy (anonymous):

Hmm.. ALright.. I am going to try to graph this.

OpenStudy (anonymous):

Alright! I got it. Thank you so much(:

OpenStudy (anonymous):

once you do, you should see that the corners are \((30,50)\) and \((20,80)\)

OpenStudy (anonymous):

then plug those numbers in to \[10x+15y\] and see which one is the biggest

OpenStudy (anonymous):

That is what I got! for the points.

OpenStudy (anonymous):

that was your original question, which was not clear but is now compute \[10\times 30+15\times 50\] and \[10\times 20+15\times 80\] to finish

OpenStudy (anonymous):

which ever is biggest is the point you want, i.e. how many of each to produce

OpenStudy (anonymous):

THANK YOU SO MUCH!!!!

OpenStudy (anonymous):

you are quite welcome hope it is clear

OpenStudy (anonymous):

Very!(: